# Quantum Algorithms in a Superposition of Spacetimes

### v2

@@ 7 $\mathsf{GapCVP}_{O\left(n^{2}\right)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
$=e^{-\frac{\pi\cdot\left\lVert\mathbf{e}_{0}\right\rVert^{2}}{\tilde{B}^{2}}}\cdot e^{-\frac{\pi\cdot\left(2\cdot\left\lVert\mathbf{e}_{0}\right\rVert\cdot\sqrt{n}\cdot\tilde{B}\right)}{\tilde{B}^{2}}}$
$\geq e^{-\frac{\pi\cdot n\cdot\kappa}{c_{1}^{2}\cdot g(n)^{2}}}\cdot e^{-\frac{\pi\cdot 2\cdot\left\lVert\mathbf{e}_{0}\right\rVert\cdot\sqrt{n}}{\tilde{B}}}$
$\geq e^{-\frac{\pi\cdot n\cdot\kappa}{c_{1}^{2}\cdot g(n)^{2}}}\cdot e^{-\frac{\pi\cdot 2\cdot n\cdot\sqrt{\kappa}}{c_{1}\cdot g(n)}}\enspace.$
To summarize, in case $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$, we got that the total variation distance between the output distributions of the circuits $C^{0}$, $C^{1}$ is bounded by
$2^{-n}+\frac{c_{0}\cdot 2\cdot n}{\sqrt{\kappa}}+\sqrt{1-e^{-\frac{\pi\cdot n\cdot\kappa}{c_{1}^{2}\cdot g(n)^{2}}}\cdot e^{-\frac{\pi\cdot 2\cdot n\cdot\sqrt{\kappa}}{c_{1}\cdot g(n)}}}\enspace.$
Now, going back to how we defined $g(n)$ and $\kappa$, we took
  • $\kappa:=c_{\kappa}\cdot n^{2}$ and $g(n):=c_{g}\cdot n^{2}$, for
  • $c_{\kappa}:=c_{0}^{2}\cdot 64$, $c_{g}:=256\cdot\frac{c_{0}}{c_{1}}\cdot\pi$ (this choice of $g(n)$ and $\kappa$ that depends on $c_{0}$, $c_{1}$ is legal, as $c_{0}$, $c_{1}$ are both positive absolute constants in $\mathbb{R}_{>0}$).
It follows that for these parameter choices we get,
$2^{-n}+\frac{c_{0}\cdot 2\cdot n}{\sqrt{\kappa}}+\sqrt{1-e^{-\frac{\pi\cdot n\cdot\kappa}{c_{1}^{2}\cdot g(n)^{2}}}\cdot e^{-\frac{\pi\cdot 2\cdot n\cdot\sqrt{\kappa}}{c_{1}\cdot g(n)}}}$
$=2^{-n}+\frac{c_{0}\cdot 2\cdot n}{8\cdot c_{0}\cdot n}+\sqrt{1-e^{-\frac{\pi\cdot n^{3}\cdot 64\cdot c_{0}^{2}}{c_{1}^{2}\cdot g(n)^{2}}}\cdot e^{-\frac{\pi\cdot 2\cdot n^{2}\cdot 8\cdot c_{0}}{c_{1}\cdot g(n)}}}$
$=2^{-n}+\frac{1}{4}+\sqrt{1-e^{-\frac{\pi\cdot n^{3}\cdot 64\cdot c_{0}^{2}}{c_{1}^{2}\cdot g(n)^{2}}}\cdot e^{-\frac{\pi\cdot 2\cdot n^{2}\cdot 8\cdot c_{0}}{c_{1}\cdot g(n)}}}$
$=2^{-n}+\frac{1}{4}+\sqrt{1-e^{-\frac{\pi\cdot n^{3}\cdot 64\cdot c_{0}^{2}}{c_{1}^{2}\cdot\left(256\cdot\frac{c_{0}}{c_{1}}\cdot\pi\cdot n^{2}\right)^{2}}}\cdot e^{-\frac{\pi\cdot 2\cdot n^{2}\cdot 8\cdot c_{0}}{c_{1}\cdot 256\cdot\frac{c_{0}}{c_{1}}\cdot\pi\cdot n^{2}}}}$
$=2^{-n}+\frac{1}{4}+\sqrt{1-e^{-\frac{1}{2^{10}\cdot\pi\cdot n}}\cdot e^{-\frac{1}{16}}}$
(for any $\underset{(\text{for any }n\in\mathbb{N})}{\leq}\frac{1}{2}+2^{-n}\enspace.$
@@
For $a,b\in\mathbb{N}$ let $\mathcal{U}_{\left(-a,b\right)}$ the discrete uniform distribution on the set $\{-a,\cdots,-1,0,1,\cdots,b\}$. For $\beta,\kappa\in\mathbb{N}$ such that $\beta$ is an even number, let the following variables:
  • $S_{0}:=\sum_{i\in[\kappa/2]}\mathcal{U}^{\left(i\right)}_{\left(-2^{\beta},2^{\beta}-1\right)}$ the random variable which is the sum of $\kappa/2$ i.i.d. samples from $\mathcal{U}_{\left(-2^{\beta},2^{\beta}-1\right)}$.
  • $S_{1}:=\sum_{i\in[\kappa/2]}\mathcal{U}^{\left(i\right)}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$ the random variable which is the sum of $\kappa/2$ i.i.d. samples from $\mathcal{U}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$.
  • We define the total sum of the two cases $S:=S_{0}+S_{1}$.
@@
  • The expectation of $\mathcal{U}^{(i)}_{\left(-2^{\beta},2^{\beta}-1\right)}$ is $\mu_{i,0}=-\frac{1}{2}$ and the expectation of $\mathcal{U}^{(i)}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$ is $\mu_{i,1}=\frac{1}{2}$.
  • The variance of both $\mathcal{U}^{(i)}_{\left(-2^{\beta},2^{\beta}-1\right)}$, $\mathcal{U}^{(i)}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$ is the same: $\sigma_{i,b}^{2}=\frac{\left(2^{\beta+1}+1\right)^{2}-1}{12}=\frac{2^{2\beta+2}+2^{\beta+2}}{12}=\Theta\left(2^{2\beta}\right)$.
@@
  • The third moment of both $\mathcal{U}^{(i)}_{\left(-2^{\beta},2^{\beta}-1\right)}$, $\mathcal{U}^{(i)}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}$ is of the same order:
    $\gamma_{i,0}:=\frac{\mathbb{E}_{u\leftarrow\mathcal{U}_{\left(-2^{\beta},2^{\beta}-1\right)}}\left(|u-\mu_{i,0}|^{3}\right)}{\sigma^{3}}=\Theta\left(\frac{1}{\kappa\sqrt{\kappa}}\right)\enspace,$
    $\gamma_{i,1}:=\frac{\mathbb{E}_{u\leftarrow\mathcal{U}_{\left(-\left(2^{\beta}-1\right),2^{\beta}\right)}}\left(|u-\mu_{i,1}|^{3}\right)}{\sigma^{3}}=\Theta\left(\frac{1}{\kappa\sqrt{\kappa}}\right)\enspace.$
  • The expectation of $S$ is $\mu:=\sum_{i\in[\kappa/2],b\in\{0,1\}}\mu_{i,b}=0$.
  • The sum of third moments is $\gamma:=\sum_{i\in[\kappa/2],b\in\{0,1\}}\gamma_{i,b}=\Theta\left(\frac{1}{\sqrt{\kappa}}\right)$.
  • The maximal absolute value that $S$ can take is $B:=\kappa\cdot 2^{\beta}$.
  • For every $i\in[\kappa/2],b\in\{0,1\}$, the random variable $S^{(i,b)}:=S-\mathcal{U}^{\left(i,b\right)}$ is summing all $\frac{\kappa}{2}$ variables from $S_{0}$ and $S_{1}$, except variable number $i$ from $S_{b}$.
@@
$\left\lVert\mathcal{D}\left(S\right)-\mathcal{D}\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)\right\rVert_{TV}$
$\leq\frac{3}{2}\cdot\sigma\cdot\sum_{i\in[\kappa/2],b\in\{0,1\}}\left(\left(\gamma_{i,b}+\frac{2\cdot\sigma_{i,b}^{2}}{3\cdot\sigma^{3}}\right)\cdot\left\lVert\mathcal{D}\left(S^{(i,b)}\right)-\mathcal{D}\left(S^{(i,b)}+1\right)\right\rVert_{TV}\right)$
@@
$\frac{3}{2}\cdot\sigma\cdot\sum_{i\in[\kappa/2],b\in\{0,1\}}\left(\left(\gamma_{i,b}+\frac{2\cdot\sigma_{i,b}^{2}}{3\cdot\sigma^{3}}\right)\cdot\left\lVert\mathcal{D}\left(S^{(i,b)}\right)-\mathcal{D}\left(S^{(i,b)}+1\right)\right\rVert_{TV}\right)$
$=\sum_{i\in[\kappa/2],b\in\{0,1\}}\left(\left(\frac{3}{2}\cdot\sigma\cdot\gamma_{i,b}+\frac{\sigma_{i,b}^{2}}{\sigma^{2}}\right)\cdot\left\lVert\mathcal{D}\left(S^{(i,b)}\right)-\mathcal{D}\left(S^{(i,b)}+1\right)\right\rVert_{TV}\right)$
=$\underset{(*)}{=}\kappa\cdot\Theta\left(\left(\frac{3}{2}\cdot\sigma\cdot\gamma_{1,0}+\frac{\sigma_{1,0}^{2}}{\sigma^{2}}\right)\cdot\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)$
$=\Theta\left(\kappa\cdot\frac{3}{2}\cdot\sigma\cdot\gamma_{1,0}\cdot\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)$
$+\Theta\left(\kappa\cdot\frac{\sigma_{1,0}^{2}}{\sigma^{2}}\cdot\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)$
$\leq O\left(2^{\beta}\cdot\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)+O\left(\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)$
$\leq O\left(2^{\beta}\cdot\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)\enspace,$
where the equality ()(*)( ∗ ) follows because for every $i\in[\kappa/2],b\in\{0,1\}$, $\gamma_{i,b}=\Theta\left(\gamma_{1,0}\right)$, $\sigma_{i,b}=\sigma_{1,0}$, and
$\left\lVert\mathcal{D}\left(S^{(i,b)}\right)-\mathcal{D}\left(S^{(i,b)}+1\right)\right\rVert_{TV}=\Theta\left(\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\right)\enspace.$
@@
$2^{\beta}\cdot\left\lVert\mathcal{D}\left(S^{(1,0)}\right)-\mathcal{D}\left(S^{(1,0)}+1\right)\right\rVert_{TV}\leq O\left(\frac{1}{\sqrt{\kappa}}\right)\enspace,$
@@
$\left\lVert\mathcal{D}\left(S\right)-\mathcal{D}\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)\right\rVert_{TV}\leq O\left(\frac{1}{\sqrt{\kappa}}\right)\enspace.$
To move to the desired distribution $D_{\tilde{B}}$, we use the triangle inequality and Fact 3.3, to get,
$\left\lVert\mathcal{D}\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)-\mathcal{D}\left(D_{\tilde{B}}\right)\right\rVert_{TV}$
$\leq\left\lVert\mathcal{D}\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)-\mathcal{D}\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\tilde{B}\right)\right)\right\rVert_{TV}$
$+\left\lVert\mathcal{D}\left(\mathcal{N}^{\kappa\cdot 2^{\beta}}\left(0,\tilde{B}\right)\right)-\mathcal{D}\left(\mathcal{N}^{\tilde{B}}\left(0,\tilde{B}\right)\right)\right\rVert_{TV}$
$+\left\lVert\mathcal{D}\left(\mathcal{N}^{\tilde{B}}\left(0,\tilde{B}\right)\right)-\mathcal{D}\left(D_{\tilde{B}}\right)\right\rVert_{TV}\leq O\left(\frac{1}{\sqrt{\kappa}}\right)\enspace.$
@@
Let $S$ and $S^{\left(i,b\right)}$ as defined in the statement and proof of Lemma 7.1. Then,
$\left\lVert\left(S^{\left(i,b\right)}\right)-\left(S^{\left(i,b\right)}+1\right)\right\rVert_{TV}\leq O\left(\frac{1}{2^{\beta}\cdot\sqrt{\kappa}}\right)\enspace.$

### v3

@@ 7 $\mathsf{GapCVP}_{O\left(n\sqrt{n}\right)}\leq_{p}\mathsf{SISD}_{\mathsf{poly},O(\log(n))}$
$=e^{-\frac{\pi\cdot\left\lVert\mathbf{e}_{0}\right\rVert^{2}}{\tilde{B}^{2}}}\cdot e^{-\frac{\pi\cdot 2\cdot\left\lVert\mathbf{e}_{0}\right\rVert\cdot\sqrt{n}}{\tilde{B}}}$
$>e^{-\frac{32^{2}\cdot n\cdot\pi\cdot d^{2}}{g(n)^{2}\cdot d^{2}}}\cdot e^{-\frac{32\cdot\sqrt{n}\cdot\pi\cdot 2\cdot d\cdot\sqrt{n}}{g(n)\cdot d}}$
$=e^{-\frac{32^{2}\cdot n\cdot\pi}{c_{g}^{2}\cdot n^{3}}}\cdot e^{-\frac{32\cdot\sqrt{n}\cdot\pi\cdot 2\cdot\sqrt{n}}{c_{g}\cdot n\sqrt{n}}}$
$=e^{-\frac{32^{2}\cdot\pi}{c_{g}^{2}\cdot n^{2}}}\cdot e^{-\frac{32\cdot\pi\cdot 2}{c_{g}\sqrt{n}}}\enspace,$
which implies
$\left(e^{-\frac{\pi\cdot\left(\left\lVert\mathbf{e}_{0}\right\rVert^{2}+2\cdot\left\lVert\mathbf{e}_{0}\right\rVert\cdot\sqrt{n}\cdot\tilde{B}\right)}{\tilde{B}^{2}}}\right)\geq\left(e^{-\frac{1}{\Omega\left(\sqrt{n}\right)}}\right)\enspace.$
Finally, note that for every positive constant $c\in\mathbb{R}_{>0}$ we have the limit
$\lim_{n\rightarrow\infty}\frac{\left(1-e^{-\frac{c}{\sqrt{n}}}\right)}{\frac{1}{\sqrt{n}}}=c\enspace.$
It follows that
$\sqrt{1-e^{-\frac{\pi\cdot\left(\left\lVert\mathbf{e}_{0}\right\rVert^{2}+2\cdot\left\lVert\mathbf{e}_{0}\right\rVert\cdot\sqrt{n}\cdot\tilde{B}\right)}{\tilde{B}^{2}}}}\underset{(\exists c\in\mathbb{R}_{>0})}{\leq}\sqrt{1-e^{-\frac{c}{\sqrt{n}}}}\leq\frac{\sqrt{c}}{n^{\frac{1}{4}}}\enspace.$
To summarize, in case $\left(\mathbf{B},\mathbf{t},d\right)\in\prod_{\text{YES}}$, we have the upper bound
$\left\lVert\mathcal{D}\left(C^{0}\right)-\mathcal{D}\left(C^{1}\right)\right\rVert_{TV}\leq\frac{1}{4}+2^{-n}+\frac{\sqrt{c}}{n^{\frac{1}{4}}}\leq\frac{1}{4}+O\left(n^{-\frac{1}{4}}\right)\enspace.$
@@
For $\kappa\in\mathbb{N}$, $\beta\in\mathbb{N}\cup\{0\}$ such that $\kappa$ is an even number, let $S$ the random variable defined as the sum of $\kappa+1$ random variables: $\kappa$ i.i.d. samples from $\mathcal{U}_{\left(0,2^{\beta+1}-1\right)}$ (the discrete uniform distribution on $\{0,1,\cdots,2^{\beta+1}-1\}$), followed by subtracting (the zero-entropy ”variable”) $\frac{\kappa}{2}\cdot\left(2^{\beta+1}-1\right)$:
$S:=\sum_{i\in[\kappa]}\mathcal{U}^{\left(i\right)}_{\left(0,2^{\beta+1}-1\right)}-\frac{\kappa}{2}\cdot\left(2^{\beta+1}-1\right)\enspace.$
@@
  • The expectation of $\mathcal{U}^{(i)}_{\left(0,2^{\beta+1}-1\right)}$ is $\mu_{i}=\frac{2^{\beta+1}-1}{2}$.
  • The variance of $\mathcal{U}^{(i)}_{\left(0,2^{\beta+1}-1\right)}$ is: $\sigma_{i}^{2}=\frac{\left(2^{\beta+1}\right)^{2}-1}{12}=\frac{2^{2\beta+2}-1}{12}=\Theta\left(2^{2\beta}\right)$.
@@
  • The third moment of $\mathcal{U}^{(i)}_{\left(0,2^{\beta+1}-1\right)}$ is of the order:
    $\gamma_{i}:=\frac{\mathbb{E}_{u\leftarrow\mathcal{U}_{\left(0,2^{\beta+1}-1\right)}}\left(|u-\mu_{i}|^{3}\right)}{\sigma^{3}}=\Theta\left(\frac{1}{\kappa\sqrt{\kappa}}\right)\enspace.$
  • The expectation of the sum of the $\kappa$ variables is $\frac{\kappa}{2}\cdot\left(2^{\beta+1}-1\right)$, and since in the full sum $S$ we subtracted it, we nullified the expectation of $S$, i.e., $\mathbb{E}\left(S\right):=\mu:=\sum_{i\in[\kappa]}\mu_{i}-\frac{\kappa}{2}\cdot\left(2^{\beta+1}-1\right)=0$.
  • The sum of third moments is $\gamma:=\sum_{i\in[\kappa]}\gamma_{i}=\Theta\left(\frac{1}{\sqrt{\kappa}}\right)$.
  • For every $i\in[\kappa]$, the random variable $S^{(i)}:=S-\mathcal{U}^{\left(i\right)}_{\left(0,2^{\beta+1}-1\right)}$ is summing all $\kappa$ variables, except variable $i$.
@@
$\left\lVert\mathcal{D}\left(S\right)-\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)\right\rVert_{TV}$
$\leq\frac{3}{2}\cdot\sigma\cdot\sum_{i\in[\kappa]}\left(\left(\gamma_{i}+\frac{2\cdot\sigma_{i}^{2}}{3\cdot\sigma^{3}}\right)\cdot\left\lVert\mathcal{D}\left(S^{(i)}\right)-\mathcal{D}\left(S^{(i)}+1\right)\right\rVert_{TV}\right)$
@@
$\frac{3}{2}\cdot\sigma\cdot\sum_{i\in[\kappa]}\left(\left(\gamma_{i}+\frac{2\cdot\sigma_{i}^{2}}{3\cdot\sigma^{3}}\right)\cdot\left\lVert\mathcal{D}\left(S^{(i)}\right)-\mathcal{D}\left(S^{(i)}+1\right)\right\rVert_{TV}\right)$
$=\sum_{i\in[\kappa]}\left(\left(\frac{3}{2}\cdot\sigma\cdot\gamma_{i}+\frac{\sigma_{i}^{2}}{\sigma^{2}}\right)\cdot\left\lVert\mathcal{D}\left(S^{(i)}\right)-\mathcal{D}\left(S^{(i)}+1\right)\right\rVert_{TV}\right)$
=$\underset{(*)}{=}\kappa\cdot\left(\left(\frac{3}{2}\cdot\sigma\cdot\gamma_{1}+\frac{\sigma_{1}^{2}}{\sigma^{2}}\right)\cdot\left\lVert\mathcal{D}\left(S^{(1)}\right)-\mathcal{D}\left(S^{(1)}+1\right)\right\rVert_{TV}\right)$
$=\kappa\cdot\frac{3}{2}\cdot\sigma\cdot\gamma_{1}\cdot\left\lVert\mathcal{D}\left(S^{(1)}\right)-\mathcal{D}\left(S^{(1)}+1\right)\right\rVert_{TV}$
$+\kappa\cdot\frac{\sigma_{1}^{2}}{\sigma^{2}}\cdot\left\lVert\mathcal{D}\left(S^{(1)}\right)-\mathcal{D}\left(S^{(1)}+1\right)\right\rVert_{TV}$
$\leq O\left(2^{\beta}\cdot\left\lVert\mathcal{D}\left(S^{(1)}\right)-\mathcal{D}\left(S^{(1)}+1\right)\right\rVert_{TV}\right)+O\left(\left\lVert\mathcal{D}\left(S^{(1)}\right)-\mathcal{D}\left(S^{(1)}+1\right)\right\rVert_{TV}\right)$
$\leq O\left(2^{\beta}\cdot\left\lVert\mathcal{D}\left(S^{(1)}\right)-\mathcal{D}\left(S^{(1)}+1\right)\right\rVert_{TV}\right)\enspace,$
where the equality ()(*)( ∗ ) follows because for every $i\in[\kappa]$, $\gamma_{i}=\gamma_{1}$, $\sigma_{i}=\sigma_{1}$, and
$\left\lVert\mathcal{D}\left(S^{(i)}\right)-\mathcal{D}\left(S^{(i)}+1\right)\right\rVert_{TV}=\left\lVert\mathcal{D}\left(S^{(1)}\right)-\mathcal{D}\left(S^{(1)}+1\right)\right\rVert_{TV}\enspace.$
@@
$2^{\beta}\cdot\left\lVert\mathcal{D}\left(S^{(1)}\right)-\mathcal{D}\left(S^{(1)}+1\right)\right\rVert_{TV}\leq O\left(\frac{1}{\sqrt{\kappa}}\right)\enspace,$
@@
$\left\lVert\mathcal{D}\left(S\right)-\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)\right\rVert_{TV}\leq O\left(\frac{1}{\sqrt{\kappa}}\right)\enspace.$
Note that in order to finish our proof, by the triangle inequality, it will be sufficient to give an upper bound of $O\left(\frac{1}{\sqrt{\kappa}}\right)$ on $\left\lVert\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)-\mathcal{D}\left(D_{\tilde{B}}\right)\right\rVert_{TV}$.
Using the triangle inequality twice we get,
$\left\lVert\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)-\mathcal{D}\left(D_{\tilde{B}}\right)\right\rVert_{TV}$
$\leq\left\lVert\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)-\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,\tilde{B}\right)\right)\right\rVert_{TV}$
$+\left\lVert\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,\tilde{B}\right)\right)-\mathcal{D}\left(\mathcal{N}^{\tilde{B}}\left(0,\tilde{B}\right)\right)\right\rVert_{TV}$
$+\left\lVert\mathcal{D}\left(\mathcal{N}^{\tilde{B}}\left(0,\tilde{B}\right)\right)-\mathcal{D}\left(D_{\tilde{B}}\right)\right\rVert_{TV}\enspace.$
Since $\tilde{B}:=\lceil\sqrt{\kappa}\cdot\tilde{\sigma}\rceil$ then $\tilde{B}-1<\sqrt{\kappa}\cdot\tilde{\sigma}\leq\tilde{B}$, and thus
$\left\lVert\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,\sqrt{\kappa}\cdot\tilde{\sigma}\right)\right)-\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,\tilde{B}\right)\right)\right\rVert_{TV}<\left\lVert\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,\tilde{B}-1\right)\right)-\mathcal{D}\left(\mathcal{N}^{\mathbb{Z}}\left(0,\tilde{B}\right)\right)\right\rVert_{TV}\enspace.$
Finally, by Fact 3.3, all of the above 3 summands are bounded by $O\left(\frac{1}{\tilde{B}}\right)\leq O\left(\frac{1}{\sqrt{\kappa}}\right)$, as needed. ∎
@@
Let $S$ and $S^{\left(1\right)}$ as defined in the statement and proof of Lemma 7.1. Then,
$\left\lVert\mathcal{D}\left(S^{\left(1\right)}\right)-\mathcal{D}\left(S^{\left(1\right)}+1\right)\right\rVert_{TV}\leq O\left(\frac{1}{2^{\beta}\cdot\sqrt{\kappa}}\right)\enspace.$